OpenStudy (wwe123):

if α,β are the roots of ax2 +bx+c=0 find :-

OpenStudy (wwe123):

\[(i) [\frac{ \alpha }{ \beta } -\frac{ \beta }{\alpha}]^{2}\]

OpenStudy (anonymous):

\[ [\frac{ \alpha }{ \beta } -\frac{ \beta }{\alpha}]^{2}=(\frac{\alpha^2 -\beta^2}{\alpha \beta})^2=\frac{\alpha^4+\beta^4-2\alpha^2 \beta^2}{\alpha^2 \beta^2}\]and note that\[\alpha^4+\beta^4=(\alpha^2+\beta^2)^2-2\alpha^2 \beta^2=((\alpha+\beta)^2-2\alpha \beta)^2-2\alpha^2 \beta^2\]

OpenStudy (wwe123):

i got my error.

OpenStudy (anonymous):

:)